OpenStudy (anonymous):

Find the nth term of the arithmetic sequence where a1 = 5000 and d = -100

OpenStudy (amistre64):

do you recall the general formula for an arithmetic seq?

OpenStudy (anonymous):

an = dn + c

OpenStudy (amistre64):

good, lets use that :)

OpenStudy (amistre64):

\[a_n=a_1+d(n-1)\] might be more applicable

OpenStudy (amistre64):

\[a_{n}=5000-100(n-1)\]

OpenStudy (amistre64):

when n=1, the first term; we get 5000

OpenStudy (amistre64):

do you know if we start ar a1 or a0 tho?

OpenStudy (anonymous):

so wouldn't the first term equal 4900?

OpenStudy (amistre64):

whats our first term? a0 or a1 ?

OpenStudy (anonymous):

a1 i guess

OpenStudy (amistre64):

then a1 = 5000 :) a2 = 4900, a3 = 4800 etc .....

OpenStudy (anonymous):

but what would be the nth term?

OpenStudy (amistre64):

\[a_{n}=5000-100(n-1)\] whatever n value you want to plug in; that is what it will be

OpenStudy (amistre64):

the nth terms IS: 5000 -100(n-1)

OpenStudy (amistre64):

you have to realize that "nth" is not static, it can take on any value it wants. When you determine you want the 253rd term, then n=253. When you want the 15th term, then n=15; etc ....